OpenStudy (anonymous):

Simplify: SqRt45+Sqrt20-SqRt5

OpenStudy (johnweldon1993):

*hint.... \[\large \sqrt{45} + \sqrt{20} - \sqrt{5}\] can be re-written as ... \[\large \sqrt{9}\sqrt{5} + \sqrt{4}\sqrt{5} - \sqrt{5}\] And since we already know that the square root of 9 is 3...and the square root of 4 is 2... \[\large 3\sqrt{5} + 2\sqrt{5} - \sqrt{5}\] Can you simplify this now??

OpenStudy (anonymous):

No Please show me the answer